Seasoned travelers know that an electrical device that works in one country may not work in...

  1. Seasoned travelers know that an electrical device that works in one country may not work in another. Imagine taking a 1500 W hair dryer made in the United States, where the rms wall voltage is 120 V, to Australia, where the rms wall voltage is 230 V. Find the power output of the hair dryer when connected to a wall outlet in Australia. (5500W)
Answer #1

as we know

P = V^2/ R

R = V^2 / P

so

P2 / P1 = (V2/V1) ^2

P2 = 1500 * ( 230 / 120)^2

P2 = 5510 W

======

Comment before rate in case any doubt, will reply for sure.. goodluck